# ReZero: Boosting MCTS-based Algorithms by Just-in-Time and Speedy Reanalyze

### v2

@@ Abstract
MCTS-based algorithms, such as MuZero and its derivatives, have achieved widespread success in various decision-making domains. These algorithms employ the reanalyze process to enhance sample efficiency, albeit at the expense of significant wall-clock time consumption. To address this issue, we propose a general approach named ReZero to boost MCTS-based algorithms. Specifically, we propose a new scheme that simplifies data collecting and reanalyzing, which significantly reduces the search cost while guarantees the performance as well. Furthermore, to accelerate each search process, we conceive a method to reuse the subsequent information in the trajectory. The corresponding analysis conducted on the bandit model also provides auxiliary theoretical substantiation for our design. Experiments conducted on Atari environments and board games demonstrates that ReZero substantially improves training speed while maintaining high sample efficiency. The code is available as part of the LightZero benchmark at https://github.com/opendilab/LightZero.
@@ 1 Introduction
As a pivotal subset of artificial intelligence, Reinforcement Learning (RL) (Sutton & Barto, 1988) has acquired achievements and applications across diverse fields, including interactive gaming (Vinyals et al., 2019), autonomous vehicles (Li et al., 2022), and natural language processing (Rafailov et al., 2023). Despite its successes, a fundamental challenge plaguing traditional model-free RL algorithms is their low sample efficiency. These algorithms typically require a large amount of data to learn effectively, making them often infeasible in the real world. In response to this problem, numerous model-based RL(Janner et al., 2019) (Hafner et al., 2020) methods have emerged. These methods involve learning an additional model of the environment from data and utilizing the model to assist the agent’s learning, thereby improving sample efficiency significantly.
@@
Within this field, Monte Carlo Tree Search (Świechowski et al., 2023) (MCTS) has been proven to be a efficient method for utilizing models for planning. It incorporates the UCB1 algorithm (Auer et al., 2002) into the tree search process and has achieved promising results in in a wide range of application scenarios. Specifically, AlphaZero (Silver et al., 2017) plays a big role in combining deep RL with MCTS, achieving notable accomplishments that can beat top-level human players. While AlphaZero can only be applied to environments with perfect simulators, MuZero (Schrittwieser et al., 2019) extended the algorithm to cases without known environment models, resulting in good performances in a wider range of tasks. Following MuZero, many successor algorithms have emerged, enabling MuZero to be applied in continuous action spaces (Hubert et al., 2021), environments with stochastic dynamics (Antonoglou et al., 2021), offline RL training scenarios (Schrittwieser et al., 2021), and etc. All these MCTS-based algorithms (Silver et al., 2016, 2017; Schrittwieser et al., 2019) made valuable contributions to the universal applicability of the MCTS+RL framework.
@@
However, the extensive tree search computations incur additional time overhead for these algorithms: During the data collection phase, the agent needs to execute MCTS to select an action every time it receives a new state. Furthermore, due to the characteristics of tree search, it is challenging to parallelize it using commonly used vectorized environments methods, further amplifying the speed disadvantage. On the other hand, during the reanalyze (Schrittwieser et al., 2021) process, in order to obtain higher-quality update targets, the latest model parameters are used to re-run MCTS on samples in buffer. The time required to process samples thus increases as a trade-off for high sample efficiency. The excessive time cost has become a bottleneck hindering the further promotion of these algorithms.
@@
Recently, a segment of research endeavors are directed toward mitigating the above wall-clock time overhead about MCTS-based algorithms. On the one hand, SpeedyZero (Mei et al., 2023) greatly diminishes algorithms’ time overheads by deploying a parallel system; however, it demands more computational resources. On the other hand, it remains imperative to identify methodologies that accelerate these algorithms without imposing extra computational demands. PTSAZero (Fu et al., 2023), for instance, compresses the search space via state abstraction, decreasing the time cost per search. Nevertheless, it is widely acknowledged that MCTS-based algorithms necessitate search times to assure their performance. From this perspective, our research try to discover strategies that preserve algorithmic performance while minimizing the search times as much as possible. Moreover, given the observation that the tree search trajectories for states at adjacent time-steps frequently exhibit a significant degree of overlap, it becomes feasible to leverage the reuse of search information across different time-steps as a method to further expedite each search process.
In this paper, we introduce a new approach designed to boost the MCTS-based algorithms. Firstly, drawing inspiration from the utilization of a periodically updated target network in DQN (Mnih et al., 2013), we find that it is not essential to reanalyze the mini-batch at each iteration. Therefore, we propose a simplified framework which periodically reanalyzes the buffer to update policy targets in a just-in-time manner. Such an approach markedly diminishes the number of calls to MCTS. Secondly, within the reanalyze process, given the collected game trajectory, we propose to reuse search information from subsequent state to accelerate the search of the current state. In addition, we provide theoretical analysis on the potential implications of these modifications about the quality of target policy. Our approach can be easily applied to the MCTS-based algorithm family and can be synergistically implemented alongside existing acceleration methods. In addition, it will not bring about an increase in hardware overhead. Empirical experiments demonstrate that our approach yields good results in both single-agent environment (Atari (Bellemare et al., 2013)) and two-player board games (Silver et al., 2016), greatly improving the training speed while maintaining high sample efficiency. Ablation experiments explore the impact of different reanalyze periods and the acceleration effect of information reuse on a single search. The main contributions of this paper can be summarized as follows:
  • We propose a simplified framework that greatly reduces the number of calls to MCTS, thus speeding up MCTS-based algorithms.
@@
  • We propose a method to speed up a single search by reusing search information. Theoretical support for the effectiveness of the method is also provided.
@@
  • We conduct serveal experiments on diverse environments and further investigate our method through ablations.
@@ 2.1 MCTS-based Algorithms
AlphaGo (Silver et al., 2016) and AlphaZero (Silver et al., 2017) combined MCTS with deep RL, achieving significant results in board games, defeating world-champion players, and revealing their remarkable capabilities. Following these achievements, MuZero (Schrittwieser et al., 2019) combines tree search with a learned value equivalent model (Grimm et al., 2020b), successfully promoting the algorithm’s application to scenarios without known models, such as Atari. Subsequent to MuZero’s development, several new works based on the MuZero framework emerged. Ye et al. (2021) further improved MuZero’s sample efficiency, using very little data for training and still achieving outstanding performance. Hubert et al. (2021) and Antonoglou et al. (2021) extended the MuZero agent to environments with large action spaces and stochastic environments, further broadening the algorithm’s usage scenario. Schrittwieser et al. (2021) applied MuZero’s reanalyze operation as a policy enhancement operator and extended it to offline settings. Danihelka et al. (2022) changed the node selection method during tree search, so the policy obtained from the search shows a monotonic improvement, ensuring algorithm performance even with low simulation numbers. Despite the aforementioned excellent improvement techniques, the current time overhead of MCTS-based RL remains significant, which is the main problem this paper aims to address.
@@ 2.2 MCTS Acceleration
Recent research has focused on accelerating MCTS-based algorithms. The work by Mei et al. (2023) significantly reduces the algorithm’s time overhead by designing a parallel system; however, this approach requires more computational resources and involves some adjustments for large batch training. Fu et al. (2023) narrows the search space through state abstraction, which amalgamates redundant spatial information, thereby reducing the time cost per search. Silver et al. (2016); Wu (2019) use information reuse to speed up the search process. They save sub-trees of the search tree and serve as initialization for the next search. However, this naive front-to-back information reuse is significantly different from our approach. To our knowledge, we are the first to enhance MCTS by reusing information in a backwards manner. Our proposed approach can seamlessly integrate with various MCTS-based algorithms, many of which (Danihelka et al., 2022; Mei et al., 2023; Fu et al., 2023; Leurent & Maillard, 2020) are orthogonal to our contributions.
@@ 3.1 MuZero
MuZero (Schrittwieser et al., 2019) is a fundamental model-based RL algorithm that incorporates a value-equivalent model (Grimm et al., 2020a) and leverages Monte Carlo Tree Search for planning within a learned latent space. The model consists of three core components: a representation model $h_{\theta}$, a dynamics model $g_{\theta}$, and a prediction model $f_{\theta}$:
@@
During the inference phase, the representation model transforms a sequence of the last $l$ observations $o_{t-l:t}$ into a corresponding latent state representation $s_{t}$. The dynamics model processes this latent state alongside an action $a_{t}$, yielding the subsequent latent state $s_{t+1}$ and an estimated reward $r_{t}$. Finally, the prediction model accepts a latent state and produces both the predicted policy $p_{t}$ and the state’s value estimate $v_{t}$. These outputs are instrumental in guiding the agent’s action selection process throughout its MCTS. Lastly the agent selects or samples the best action $a_{t}$ following the searched visit count distribution. During the training phase, given a training sequence $\{o_{t-l:t+K},a_{t+1:t+K},u_{t+1:t+K},\pi_{t+1:t+K},z_{t+1:t+K}\}$ at time $t$ sampled from the replay buffer, where $u_{t+k}$ denotes the actual reward obtained from the environment, $\pi_{t+k}$ represents the target policy obtained through MCTS during the agent-environment interaction, and $z_{t+k}$ is the value target computed using n-step bootstrapping (Hessel et al., 2018). The representation model initially converts the sequence of observations $o_{t-l:t}$ into the latent state $s_{t}^{0}$. Subsequently, the dynamic model executes $K$ latent space rollouts based on the sequence of actions $a_{t+1:t+K}$. The latent state derived after the $k$-th rollout is denoted as $s_{t}^{k}$, with the corresponding predicted reward indicated as $r_{t}^{k}$. Upon receiving $s_{t}^{k}$, the prediction model generates a predicted policy $p_{t}^{k}$ and a estimated value $v_{t}^{k}$. The final training loss encompasses three components: the policy loss ($l_{p}$), the value loss ($l_{v}$), and the reward loss ($l_{r}$):
$L_{\text{MuZero}}=$ $\sum_{k=0}^{K}l_{p}(\pi_{t+k},p_{t}^{k})+\sum_{k=0}^{K}l_{v}(z_{t+k},v_{t}^{k})+\sum_{k=1}^{K}l_{r}(u_{t+k},r_{t}^{k})$ (2)
MuZero Reanalyze, as introduced in (Schrittwieser et al., 2021), is an advanced iteration of the original MuZero algorithm. This variant enhances the model’s accuracy by conducting a fresh Monte Carlo Tree Search on sampled states with the most recent version of the model, subsequently utilizing the refined policy from this search to update the policy targets. Such reanalysis yields targets of superior quality compared to those obtained during the initial data collection phase. The Schrittwieser et al. (2021) expands upon this approach, formalizing it as a standalone method for policy refinement. This innovation opens avenues for its application in offline settings, where interactions with the environment are not possible. Traditional uses of the algorithm have intertwined reanalysis with training, periodically revisiting a subset of samples within a mini-batch for reevaluation. Furthermore, we suggests a paradigm shift, advocating for the decoupling of the reanalyze process from training routines, thus providing a more flexible and potentially more efficient methodology.
@@ 3.2 Stochastic Bandits
In the upcoming Section 4.2.2, we plan to offer an in-depth look at the theoretical assurances behind the performance of our newly designed ReZero agent, which is rooted in the fundamentals of bandit theory. Before diving into that, we’ll start with an easy-to-understand overview of the key concepts surrounding Stochastic Bandits.
A stochastic bandit $\nu$ with $K$ arms (Simchi-Levi et al., 2023) can be formulated as $\nu=\{P_{i}|i\in{1,2,...,K}\}$, where $P_{i}$ denotes the reward distribution corresponding to the arm $i$. The expected reward of arm $i$ is denoted as $\mu_{i}(\nu)$. Given an algorithm $\pi$ interacting with the bandit $\nu$ for $n$ rounds, we define the regret $R_{n}$ as follows:
@@
where $\mu^{*}(\nu)=\max_{a\in[K]}\mu_{a}(\nu)$, $X_{t}$ denotes the random reward achieved at time $t$.
The regret quantifies the loss of reward incurred by $\pi$ compared to the optimal policy on $\nu$ during $n$ rounds of interactions. In practical problems, we aim for $\pi$ to exhibit sub-linear regret growth across a category of bandits $\Omega$ , i.e.:
@@
where sub-linear growth rate indicates $\pi$ is choosing the optimal arm almost all of the time with $n$ tends to infinity. UCB (Auer et al., 2002) is a classical bandit algorithm that achieves regret growth rate of $O(\log(n))$.
AlphaZero (Silver et al., 2017) considers the child nodes in tree search as bandit arms and applies a variant of the UCB in the tree search process:
@@ 4 Method
Figure 1: Execution workflow and runtime cycle graph about MuZero and ReZero in both single and multiple worker cases. The number inside the modules represent the number of iterations, and the number under the modules represent the time required for module execution. The model is updated $n$ iterations between two collections. MuZero reanalyzes the mini-batch before each model update. ReZero reanalyzes the entire buffer after certain update iterations($\frac{n}{3}$ for example), which not only reduces the total number of MCTS calls, but also takes advantage of the processing speed of large batches.
In this section, we introduce two techniques to boost MCTS-based algorithms. Section 4.1 describes the design of just-in-time reanalyze, which achieves sufficient training on policy target meanwhile significantly reduces the number of MCTS executions. Furthermore, section 4.2 introduces speedy reanalyze: an acceleration method based on information reuse, which utilizes search results from the subsequent time-step to speed up the search at the previous time-step, thereby saving the time required for a single MCTS execution.
4.1 Just-in-Time Reanalyze: A Simplified Framework
Reanalyze, as a method to enhance the quality of training targets, was previously bound together with the training process in earlier implementations (Schrittwieser et al., 2021) (Figure 1). However, recomputing the target policy at each training iteration incurs massive MCTS computations. This dilemma motivates us to find a method to reduce the amount of MCTS calculations while ensuring performance.
We begin our detailed analysis with the following perspective: consider the reanalyze process as a policy improvement operator $\mathcal{R}_{\theta^{t}}$, which takes the current policy $P_{\theta^{t}}$ and transforms it into the target policy $T_{\theta^{t}}$ for training.
$T_{\theta^{t}}(a|s)=(\mathcal{R}_{\theta^{t}}P_{\theta^{t}})(a|s)=\frac{E_{T_{a}\sim\theta^{t}}[T_{a}]}{N}$ (6)
where $N$ is the number of simulations during the tree search and $E[T_{a}]$ is the expected total number that action $a$ was chosen. During the training process, we update the model parameter $\theta$ according to the following equation:
$\theta^{t+1}=\arg\min_{\theta}E_{s\sim\mathcal{D}}[\sum_{k=0}^{K}l^{p}(T_{\theta^{t}}(\cdot|s^{k})),P_{\theta}(\cdot|s^{k}))+l_{reg}]$ (7)
where $s$ is the state sampled from buffer $\mathcal{D}$, and $s^{k}$ is the state after $k$ unroll steps. We consider $l_{r}$ and $l_{v}$ in Equation 2 as a regularizer for $l_{p}$, and mark them as $l_{reg}$. In this way, the training can be seen as a projection operator $\mathcal{P}$ (Grill et al., 2020), projecting $T_{\theta^{t}}$ onto the policy space. The alternation of the reanalyze process and the training process constitutes policy iteration (Sutton & Barto, 1988), which can be seen as the composition of two operators:
$P_{\theta^{t+1}}=(\mathcal{P}\circ\mathcal{R}_{\theta^{t}})P_{\theta^{t}}$ (8)
In previous work (Schrittwieser et al., 2021), the reanalyze process is executed on the mini-batch of each training iteration. Changing the projection target frequently may result in an insufficient optimization, thereby degrading the quality of $\mathcal{P}$. In this way, even if the improvement operator $\mathcal{R}_{\theta}$ can produce high-quality training targets, the composite operator may not perform well. Therefore, we can adopt the practice of periodically updating the target network in DQN (Mnih et al., 2013), using Just-in-Time Reanalyze to change the projection targets only after several rounds of training.
Specifically, we propose the periodical reanalyze mechanism, i.e., we reanalyze the whole buffer after a fixed number of iterations. For each training iteration, we only need to sample the mini-batch and execute the gradient descent. Compared with previous methods, our method greatly reduces the number of MCTS, and can improve the parallelism due to the bigger search batch (Keskar et al., 2016), further enhanced the overall speed. Since the reanalyze process provides the target policy, the tree search during the collection phase is no longer necessary in the entire pipeline. Thus, we can bypass the tree search and directly select actions based on the policy, further reducing MCTS executions.
Based on our simplified framework, there are several potential advantages that may be explored in future work:
  • When directly using policy for data collection, action selection is no longer constrained by the tree search. As a result, previous vectorized environments such as (Weng et al., 2022) can be integrated to speed up. Besides, this design also makes MCTS-based algorithms compatible to existing RL exploration methods like (Badia et al., 2020).
4.2 Speedy Reanalyze: Temporal Information Reuse
The above framework reduces the total number of MCTS. Next, we propose an temporal information reuse method to speed up the single tree search (Hereafter, termed information reuse for brevity). Section 4.2.1 introduces motivations and specific details. Section 4.2.2 carries out theoretical analysis of the method based on the bandit model.
4.2.1 Reuse the Information from Backward
As a local planning approach (Świechowski et al., 2023), MCTS is executed when receives each environment’s state. The agent, by performing tree search on the current state, can gain the root value and the visit counts: the former tells the quality of the state, and the later tells which actions are more likely to yield good returns. These can be seen as a kind of information. The trajectories traversed by the search of adjacent time-steps often have many overlaps, indicating that it involves similar knowledge. If the information gained at one time-step can be passed on to another, the search process can be accelerated: for example, if the search conducted at time-step $A$ reveals the quality of state $s$, then at time-step $B$, the agent need not to spend extra search cost to explore state $s$. Therefore, we designed a method to speed up the reanalyze process by reusing search information.
Figure 2: An illustration about the information reuse between adjacent time-steps. We search in the reverse order of the trajectory. After performing search on $s_{t+1}$, we obtain the search tree $T_{t+1}$ and the root value $m_{t+1}$. When searching on $s_{t}$, we set the value of the node corresponding to $s_{t+1}$ in search tree $T_{t}$ as $m_{t+1}$. If action $a_{t}$ is selected, the traverse process will be terminated, otherwise the traverse will proceed normally.

### v3

# ReZero: Boosting MCTS-based Algorithms by Backward-view and Entire-buffer Reanalyze

@@ Abstract
Monte Carlo Tree Search (MCTS)-based algorithms, such as MuZero and its derivatives, have achieved widespread success in various decision-making domains. These algorithms employ the reanalyze process to enhance sample efficiency from stale data, albeit at the expense of significant wall-clock time consumption. To address this issue, we propose a general approach named ReZero to boost tree search operations for MCTS-based algorithms. Specifically, drawing inspiration from the one-armed bandit model, we reanalyze training samples through a backward-view reuse technique which obtains the value estimation of a certain child node in advance. To further adapt to this design, we periodically reanalyze the entire buffer instead of frequently reanalyzing the mini-batch. The synergy of these two designs can significantly reduce the search cost and meanwhile guarantee or even improve performance, simplifying both data collecting and reanalyzing. Experiments conducted on Atari environments and board games demonstrate that ReZero substantially improves training speed while maintaining high sample efficiency. The code is available as part of the LightZero benchmark at https://github.com/opendilab/LightZero.
@@ 1 Introduction
As a pivotal subset of artificial intelligence, Reinforcement Learning (RL) [1] has acquired achievements and applications across diverse fields, including interactive gaming [2], autonomous vehicles [3], and natural language processing [4]. Despite its successes, a fundamental challenge plaguing traditional model-free RL algorithms is their low sample efficiency. These algorithms typically require a large amount of data to learn effectively, making them often infeasible in the real-world scenarios. In response to this problem, numerous model-based RL [5, 6] methods have emerged. These methods involve learning an additional model of the environment from data and utilizing the model to assist the agent’s learning, thereby improving sample efficiency significantly.
@@
Within this field, Monte Carlo Tree Search [7] (MCTS) has been proven to be a efficient method for utilizing models for planning. It incorporates the UCB1 algorithm [8] into the tree search process and has achieved promising results in in a wide range of scenarios. Specifically, AlphaZero [9] plays a big role in combining deep reinforcement learning with MCTS, achieving notable accomplishments that can beat top-level human players. While it can only be applied to environments with perfect simulators, MuZero [10] extended the algorithm to cases without known environment models, resulting in good performances in a wider range of tasks. Following MuZero, many successor algorithms have emerged, enabling MuZero to be applied in continuous action spaces [11], environments with stochastic dynamics [12], offline RL training scenarios [13], and etc. All these MCTS-based algorithms [9, 10, 14] made valuable contributions to the universal applicability of the MCTS+RL paradigm.
@@
However, the extensive tree search computations incur additional time overhead for these algorithms: During the data collection phase, the agent needs to execute MCTS to select an action every time it receives a new state. Furthermore, due to the characteristics of tree search, it is challenging to parallelize it using commonly used vectorized environments [15], further amplifying the speed disadvantage. On the other hand, during the reanalyze [13] process, in order to obtain higher-quality update targets, the latest model parameters are used to re-run MCTS on the training mini-batch. The time required to process samples thus increases as a trade-off for high sample efficiency. The excessive time cost has become a bottleneck hindering the further promotion of these algorithms.
@@
Recently, a segment of research endeavors are directed toward mitigating the above wall-clock time overhead about MCTS-based algorithms. On the one hand, SpeedyZero [16] greatly diminishes algorithms’ time overheads by deploying a parallel execution training pipeline; however, it demands additional computational resources. On the other hand, it remains imperative to identify methodologies that accelerate these algorithms without imposing extra demands. PTSAZero [17], for instance, compresses the search space via state abstraction, decreasing the time cost per search. Nevertheless, it is widely acknowledged that MCTS-based algorithms necessitate search times (the number of simulations) to assure their performance. From this perspective, our research try to discover strategies that preserve algorithmic performance while minimizing the search times as much as possible.
In this paper, we introduce ReZero, a new approach designed to boost the MCTS-based algorithms. Firstly, inspired by the one-armed bandit model [18], we propose a backward-view reanalyze technique that proceeds in the reverse direction of the trajectories, utilizing previously searched root values to bypass the exploration of specific child nodes, thereby saving time. We present several propositions to articulate our insights and analyzed the convergence of the algorithm based on the bandit model. Secondly, to better adapt to our proposed backward-view reanalyze technique, we have devised a novel pipeline that concentrates MCTS calls within the reanalyze process and periodically reanalyzes the entire buffer after a fixed number of training iterations. This entire-buffer reanalyze not only reduces the number of MCTS calls but also better leverages the speed advantages of parallelization. Our approach can be easily applied to the MCTS-based algorithm family and synergistically implemented alongside existing acceleration methods. In addition, it will not bring about any overhead in computation resources. Empirical experiments (Section 5) show that our approach yields good results in both single-agent environment (Atari [19]) and two-player board games [14], greatly improving the training speed while maintaining or even improving sample efficiency. Ablation experiments explore the impact of different reanalyze frequencies and the acceleration effect of backward-view reuse on a single search. The main contributions of this paper can be summarized as follows:
@@
  • We design a method to speed up a single search by the backward-view reanalyze technique. Theoretical support for the effectiveness of our proposed method is also provided.
  • We propose a simplified framework with the entire-buffer reanalyze mechanims that further reduces the number of MCTS calls and enhance its parallelization, thus boosting MCTS-based algorithms.
@@
  • We conduct several experiments on diverse environments and investigate ReZero through ablations.
@@ 2.1 MCTS-based Algorithms
AlphaGo [14] and AlphaZero [9] combined MCTS with deep RL, achieving significant results in board games, defeating world-champion players, and revealing their remarkable capabilities. Following these achievements, MuZero [10] combines tree search with a learned value equivalent model [20], successfully promoting the algorithm’s application to scenarios without known models, such as Atari. Subsequent to MuZero’s development, several new works based on the MuZero framework emerged. EfficientZero [21] further improved MuZero’s sample efficiency, using very little data for training and still achieving outstanding performance. Another works [11, 12] extended the MuZero agent to environments with large action spaces and stochastic environments, further broadening the algorithm’s usage scenario. MuZero Unplugged [13] applied MuZero’s reanalyze operation as a policy enhancement operator and extended it to offline settings. Gumbel MuZero [22] changed the node selection method during tree search, so the policy obtained from the search shows a monotonic improvement, ensuring algorithm performance even with low simulation numbers. Despite the aforementioned excellent improvement techniques, the current time overhead of MCTS-based RL remains significant, which is the main problem this paper aims to address.
@@ 2.2 MCTS Acceleration
Recent research has focused on accelerating MCTS-based algorithms. The work [16] reduces the algorithm’s time overhead by designing a parallel system; however, this method requires more computational resources and involves some adjustments for large batch training. PTSAZero [17] narrows the search space through state abstraction, which amalgamates redundant information, thereby reducing the time cost per search. KataGo [23] use a naive information reuse trick. They save sub-trees of the search tree and serve as initialization for the next search. However, our proposed backward-view reanalyze technique is fundamentally different from this naive forward-view reuse and can enhance search results while saving time. To our knowledge, we are the first to enhance MCTS by reusing information in a backward-view. Our proposed approach can seamlessly integrate with various MCTS-based algorithms, many of which [16, 17, 22, 24] are orthogonal to our contributions.
@@ 3.1 MuZero
MuZero [10] is a fundamental model-based RL algorithm that incorporates a value-equivalent model [25] and leverages MCTS for planning within a learned latent space. The model consists of three core components: a representation model $h_{\theta}$, a dynamics model $g_{\theta}$, and a prediction model $f_{\theta}$:
@@
The representation model transforms a last observation sequences $o_{t-l:t}$ into a corresponding latent state $s_{t}$. The dynamics model processes this latent state alongside an action $a_{t}$, yielding the subsequent latent state $s_{t+1}$ and an estimated reward $r_{t}$. Finally, the prediction model accepts a latent state and produces both the predicted policy $p_{t}$ and the state’s value $v_{t}$. These outputs are instrumental in guiding the agent’s action selection throughout its MCTS. Lastly the agent samples the best action $a_{t}$ following the searched visit count distribution. MuZero Reanalyze is an advanced version of the original MuZero algorithm. This variant enhances the model’s accuracy by conducting a fresh MCTS on sampled states with the most recent version of the model, subsequently utilizing the refined policy from this search to update the policy targets. Such reanalysis yields targets of superior quality compared to those obtained during the initial data collection. Traditional uses of the algorithm have intertwined reanalysis with training, while we suggests a novel paradigm, advocating for the decoupling of the reanalyze process from training iterations, thus providing a more flexible and efficient methodology.
@@ 3.2 Stochastic Bandits
In the upcoming Section 4.1, we plan to provide an inspiration case study behind our method, which is rooted in the fundamentals of bandit algorithms. Before diving into that, we’ll start with an overview of the key concepts surrounding stochastic bandits.
A stochastic bandit $\nu$ with $K$ arms [26] can be formulated as $\nu=\{P_{i}|i\in{1,2,...,K}\}$, where $P_{i}$ denotes the reward distribution corresponding to the arm $i$. The expected reward of arm $i$ is denoted as $\mu_{i}(\nu)$. Given an algorithm $\pi$ interacting with the bandit $\nu$ for $n$ rounds, we define the regret $R_{n}$:
@@
where $\mu^{*}(\nu)=\max_{a\in[K]}\mu_{a}(\nu)$, $X_{t}$ denotes the random reward at time $t$. The regret quantifies the loss of reward incurred by $\pi$ compared to the optimal policy on $\nu$ during $n$ rounds of interactions. In practice, we aim for $\pi$ to exhibit sub-linear regret growth across a category of bandits $\Omega$ , i.e.:
@@
where sub-linear growth rate indicates $\pi$ is choosing the optimal arm almost all of the time with $n$ tends to infinity. UCB [8] is a classical bandit algorithm that achieves regret growth rate of $O(\log(n))$.
AlphaZero [9] considers the child nodes in tree search as bandit arms and applies a variant of the UCB in the tree search process:
@@ 4 Method
4.1 An Inspiring Case Study
As a classic algorithm in the realm of stochastic bandits, Upper Confidence Bound 1 (UCB1) [8] has been deployed to select nodes within a search tree.
Figure 1: The connection between MCTS and bandits. Left shows the tree search in a bandit-view. When action $A$ is selected, a reward $r_{A}$ will be returned, where $r_{A}=\sum_{t=1}^{3}\gamma^{t-1}r_{t}^{a}+\gamma^{3}v_{a}$. For the root node, the traversal, evaluation and back-propagation occurring in the gray box can be approximated as sampling from an non-stationary distribution. Thus the root node can be seen as a non-stationary bandit. Right shows the one-armed bandit case. Once the reward $\mu_{A}$ is known, we can evaluate arm A using $\mu_{A}$, thereby eliminating the need to rely on subsequent tree search processes.
As explained in Figure 1, each node in the tree can be regarded as a multi-armed bandit Notice that in the context of stochastic bandits, the reward for each arm is sampled from a stationary probability distribution. However, in the process of tree search, the value obtained from selecting a node depends on subsequent search process. The UCT [27] algorithm models this process as sampling from a non-stationary distribution which converges. For the sake of analytical convenience in this paper, we approximate the root node as a stochastic bandit and consider each child node as an arm with a deterministic reward distribution. Based on this modeling, we can analyze the Eq. 4 employed in AlphaZero [9] and derive the following proposal:
Proposal 1
Denote the prior for arm $i$ as $P_{i}$. Without loss of generality, suppose $\mu^{*}=\mu_{1}$. Assume that the reward distribution of each arm $i$ is $1$-subgaussian and $P_{1}>\mu_{1}$, then the regret $R_{n}$ of using Eq. 4 to choose arm satisfies
The assumption that $P_{1}>\mu_{1}$ is made purely for the sake of formality; in fact, Appendix A indicates that, as long as there exists a $t$ that $P_{1}\sqrt{t-1}>\mu_{1}$, a similar upper bound on regret can always be obtained. This elucidates that selecting actions according to Eq. 4 can yield a sublinear growth of regret, indicating that the final visit distribution will be concentrated on the optimal arm (see analysis for Lemma A.1 in Appendix A). This also confirms the rationality of selecting the action with the largest number of visits. The inspiration of our enhanced method comes from the one-armed bandit model: if the expected reward of a particular arm is known as prior, the necessity for exploratory trials on this arm is obviated, thereby reducing the regret upper bound. To concretize this insight, we can extend Eq. 4 and deduce the following proposition:
Proposal 2
For a bandit which satisfies the assumptions in Proposal 1, the regret upper bound can be reduced if we use the known expectation value of an arm to replace its score in Eq. 4. Specifically, if the known arm is the optimal arm with value $\mu_{1}$, then the regret $R_{n}$ satisfies:
$R_{n}$ $\leq\sum_{i:\Delta_{i}>0,i\neq l}\inf_{\varepsilon\in(0,\Delta_{i})}\Delta_{i}(3+\frac{16}{\varepsilon^{4}P_{1}^{2}}+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon})+\inf_{\varepsilon\in(0,\Delta_{l})}\Delta_{l}(1+\frac{16}{\varepsilon^{4}P_{1}^{2}})$ (7)
Proposal 2 demonstrates that, given the known expected reward of a suboptimal arm, the asymptotic upper bound of regret can be further reduced, thereby concentrating the visit distribution more heavily on the optimal arm. Meanwhile, as shown in Figure 1, when the expectation $\mu_{A}$ is known, we no longer need to evaluate arm A through real-time search results. Instead, we can directly substitute $\mu_{A}$ for the score in Eq. 4, thus saving the time spent on traverse and back-propagation processes. This inspires us to evaluate a particular child node using existing information, thereby eliminating the exploration of the corresponding sub-tree and accelerating the search process. See Appendix A for detailed proof and analysis of the two proposals.
4.2 Backward-view Reanalyze
Figure 2: An illustration about the backward-view reanalyze in both node and batch view. We sample $n+1$ trajectories of length $k+1$ to form a batch and conduct the search in the reverse direction of trajectories. From the node view, we would first search $S_{0}^{1}$ and then pass root value $m_{0}^{1}$ to $S_{0}^{0}$ to evaluate the value of a child node. $T_{0}^{1}$ and $T_{0}^{0}$ are the corresponding search trees. From the batch view, we would group all $S^{1}$s into a sub batch to search together and pass the root values to the $S^{0}$s.
Inspired by the aforementioned motivation, we aspire to obtain the reward expectation of a child node in advance to assess it. However, the true expectation is always unknown. Consequently, we resort to using the root value obtained from MCTS as an approximate substitute. Specifically, for two adjacent time steps $S_{0}^{1}$ and $S_{0}^{0}$ in Figure 2, when searching for state $S_{0}^{1}$, the root node corresponds to a child node of $S_{0}^{0}$. Therefore, the root value obtained can be utilized to assist the search for $S_{0}^{0}$. However, there is a temporal contradiction. $S_{0}^{1}$ is the successor state of $S_{0}^{0}$, yet we need to complete the search for $S_{0}^{1}$ first. Fortunately, this is possible during the reanalyze process.